### v1

# Forging Tokens for Improved Storage-efficient Training

Abstract
Recent advancements in Deep Neural Network (DNN) models have significantly improved performance across computer vision tasks. However, achieving highly generalizable and high-performing vision models requires extensive datasets, leading to large storage requirements. This storage challenge poses a critical bottleneck for scaling up vision models. Motivated by the success of discrete representations, SeiT proposes to use Vector-Quantized (VQ) feature vectors (i.e., tokens) as network inputs for vision classification. However, applying traditional data augmentations to tokens faces challenges due to input domain shift. To address this issue, we introduce TokenAdapt and ColorAdapt, simple yet effective token-based augmentation strategies. TokenAdapt realigns token embedding space for compatibility with spatial augmentations, preserving the model’s efficiency without requiring fine-tuning. Additionally, ColorAdapt addresses color-based augmentations for tokens inspired by Adaptive Instance Normalization (AdaIN). We evaluate our approach across various scenarios, including storage-efficient ImageNet-1k classification, fine-grained classification, robustness benchmarks, and ADE-20k semantic segmentation. Experimental results demonstrate consistent performance improvement in diverse experiments, validating the effectiveness of our method. Code is available at https://github.com/naver-ai/tokenadapt.
1 Introduction
Significant advances in Deep Neural Network (DNN) models [16, 48] have substantially improved their performance across various computer vision tasks. However, achieving highly generalizable and high-performing vision models demands an extensive dataset, often reaching billions of data points [25, 44, 34] with large storage requirements (e.g., the LAION-5B [42] dataset requires a 240TB storage size). This storage challenge emerges as a critical bottleneck for scaling up vision models. Consequently, optimizing storage efficiency becomes important for the practical advancement of scaled-up vision models.
Images require more storage than texts, due to their intrinsic complexity in converting continuous natural light signals into discrete pixels for computer representation. However, a considerable portion of image data corresponds to redundant or unnecessary details, such as repeated patterns or imperceptible details [39]. These properties of image data lead to large storage requirements and also induce negative impacts on visual recognition tasks as mentioned in [10, 21, 33, 1, 20, 43].
To build a deep representation away from redundant and superfluous information in image details, earlier researches [49, 17, 56, 3, 2] have explored ways to extract condensed information from images. One notable study, VQ-VAE [49], emphasized the importance of discrete representations, suggesting it as a potentially more natural fit for various modalities. They demonstrated that an image can be sufficiently expressed as a combination of vector-quantized (VQ) feature vectors by achieving similar performance as their continuous counterparts in the generation task. Recent studies [17, 39] further support this claim by successfully synthesizing high-resolution images from discrete feature vectors. Building on this understanding, SeiT [36] delved into optimizing storage efficiency by utilizing discrete representations (i.e., tokens) as network inputs. SeiT suggests using the vector-quantized feature instead of image pixels, which substantially reduces the storage size for a dataset by excluding redundancies in images.
These findings highlight the effectiveness of leveraging pre-extracted tokens. They successfully capture the essence of original pixel data while taking advantage of the inherent storage efficiency in the VQ technique. Nevertheless, this approach introduces a significant challenge: applying existing data augmentations designed for pixel images to token embedding is hindered by input domain shift. Previous studies [12, 11, 48] have highlighted the crucial role of data augmentation in model performance, positioning it as a fundamental building block extensively employed in various training recipes. Considering its significance in DNN-based models, addressing the challenges posed by the application of data augmentation to token-based vision models is crucial. This resolution holds the key to overcoming major performance bottlenecks, thereby enhancing the effectiveness of these models across diverse vision tasks, including semi-, self-supervised learning, and limited data scenarios.
To tackle this issue, we propose two simple yet effective augmentation strategies called TokenAdapt and ColorAdapt. We designed TokenAdapt motivated by the inherent property of token embedding that encapsulates image coordinate information and can be decoded back to the original image space. The purpose of our TokenAdapt is to realign the token embedding space for better compatibility with augmentation. We achieve this by converting token embedding into a learned feature space suitable for pixel-based augmentation techniques, applying the augmentations, and then reverting the augmented feature back to the original token embedding space. In this way, we can reuse established image-augmentation policies as-is. Besides, it allows us to keep the pre-trained weights of the original token-based model unchanged, avoiding significant computational overhead such as fine-tuning.
While TokenAdapt can handle geometric data augmentations, applying color-based data augmentations, which typically operate on pixel values, to a non-pixel domain remains a non-trivial task. Here, we introduce ColorAdapt, a method designed to enable controllable color changes for token embedding. Inspired by Adaptive Instance Normalization (AdaIN [24]), ColorAdapt aims to preserve object structure while introducing significant color variations. In the training phase, we employ TokenAdapt and ColorAdapt for data augmentation. With them, we find that our strategy can improve the effectiveness of token-based training in terms of model performance.
We validate the effectiveness of our proposed token-based augmentation strategies through experiments across various scenarios where data augmentation is crucial: (1) storage-efficient ImageNet-1k classification, (2) transfer learning on fine-grained classification, (3) evaluation on robustness benchmarks, and (4) ADE-20k semantic segmentation. As depicted in Figure 1, our method consistently outperforms comparison methods in storage-efficient ImageNet-1k classification with limited storage sizes. These results confirm that our proposed method significantly contributes to the storage efficient model by augmenting data effectively. Furthermore, we showcase how our augmentation strategies enhance the generalizability of trained models, as evidenced by our experimental results on fine-grained classification and robustness evaluation. Lastly, our method’s applicability to an alternative input format (e.g., DCT coefficients) demonstrates its extensibility.
2 Related Work
Learning with tokenization. The use of discrete tokens for training vision models has recently gained attention, particularly in self-supervised representation learning. BeiT [4] employs the Masked Image Modeling (MIM) framework to recover discrete tokens from masked images and CIM [18] utilizes tokens in Corrupted Image Modeling. MAGE [30] demonstrates the incorporation of tokens from VQGAN [17] for both generative training and representation learning within a unified training framework. However, existing studies mainly focus on self-supervised representation learning and rely on online tokenization, converting images to tokens during each iteration of model training. This approach incurs notable memory consumption and computational costs, limiting the full exploitation of the storage efficiency of discrete tokens.
Storage-efficient vision training. Storage-efficient vision training aims to reduce the dataset storage size. To this end, dataset distillation (DD) [53] creates a condensed dataset by leveraging the knowledge of the original dataset. Recent research [29, 40, 41, 59, 60] highlights the effectiveness of dataset distillation in improving training efficiency. However, its high computational complexity poses challenges, particularly when applied to large-scale datasets like ImageNet-1k [14]. Sampling-based methods [37, 8, 26] address this by selecting a subset of the most representative samples from the full dataset. Paul et al. [37] identify important samples very early in training. Jiang et al. [26] utilize a consistency score based on statistics collected during training. SVP [8] constructs small proxy models to find important samples. However, the diverse sample selection is not guaranteed, particularly in low data regimes, leading to suboptimal performance compared to dataset distillation (DD) methods [60]. To tackle these issues, SeiT [36] proposes a storage-efficient vision training framework by leveraging the storage efficiency of tokens. They utilize the ViT-VQGAN [56] tokenizer to extract tokens from images and store them for vision model training. As a result, they outperform other storage-efficient training methods, significantly reducing the dataset storage size.
Data augmentation. Data augmentation serves as a fundamental building block to improve the model performance across diverse computer vision tasks. Various data augmentation techniques have been employed for improved performance [15, 61] or model robustness [19, 32]. Mixup [58] introduces convex combinations of pairs of images and their labels to create a smoother decision boundary. CutMix [57] replaces certain regions of an image with a patch from another image, enhancing the model’s localization ability. AutoAugment [11] leverages reinforcement learning to automate optimal augmentation policy discovery. RandAugment [12] introduces a potent policy based on random combinations of different data augmentation methods. DeiT [48] emphasizes the crucial role of data augmentation in vision transformers for improved model performance. Previous studies extensively explore the effective data augmentation policies for pixel-based vision model training. However, we observe that pixel-based data augmentations are not well-suited for tokens, posing a critical performance bottleneck. Motivated by this, we tackle the data augmentation challenge, which has not been explored in previous studies for vision model training with tokens.
3 Method
This section starts with analyzing the attributes of tokens that make pixel-based data augmentations ineffective and unsuitable, as discussed in Section 3.1. Then, we introduce our two alternative augmentation strategies, namely TokenAdapt and ColorAdapt, designed for token-based vision model training. In Section 3.2, we present the TokenAdapt module, which facilitates the application of pixel-based data augmentations to token embedding by transforming them into a compatible domain. In Section 3.3, we propose ColorAdapt, an additional color-based token augmentation, to further diversify the training data. Finally, Section 3.4 describes the overall training procedure that incorporates our augmentation strategies for vision model training.
3.1 Limited augmentations for tokens
We examine why pixel-based geometric data augmentations are not well-suited for tokens, identifying two main reasons: (1) spatial information collapse during tokenization and (2) inter-token dependencies. The discrete tokenization process condenses an 2D image patch into a single -dimensional 1D vector (i.e., token embedding), causing a collapse of spatial information. This poses challenges for data augmentation, particularly for operations like horizontal flip (hFlip) that depend on the spatial details of the original pixels. We also observe that neighboring -dimensional token embeddings are interdependent, in contrast to the independence of individual pixels in an image. As a result, applying data augmentations related to interpolation (e.g., Resize or Mixup [58]) introduces undesired artifacts.
| Framework | hFlip | Mixup [58] | Affine | Top 1 Acc. |
| SeiT [36] | ✘ | ✘ | ✘ | 74.0 |
| ✔ | ✘ | ✘ | 73.2 (-0.8) | |
| ✔ | ✔ | ✘ | 72.0 (-2.0) | |
| ✔ | ✔ | ✔ | 71.3 (-2.7) |
Figure 2 illustrates examples related to the failure of pixel-based augmentations for tokens. When applying a hFlip, we observe the disruption of both the diagonal line (upper row) and the cat’s silhouette (lower row) due to spatial information collapse. Specifically, only the inter-token spatial relationship is flipped, while the inner-token spatial arrangement remains unchanged. We also witness that augmentations involving interpolation can lead to undesired artifacts, such as color leaking in the case of Random Resized Crop (RRC), because of interdependence between neighboring tokens. These problems are observed in the tokens extracted from natural images as well (see Figure 3). Consequently, direct application of pixel-based augmentations to token embedding negatively affects token-based vision classifier learning as shown in Table 1.
3.2 TokenAdapt
The concept of TokenAdapt is to establish a feature space that accommodates conventional pixel-based augmentations while facilitating the conversion and inversion of pre-extracted tokens under minimal computation overheads. This approach offers several advantages for our TokenAdapt. Firstly, it enables the direct reuse of well-designed pixel-based data augmentation policies. Besides, the augmented token can be easily computed via the conversion-augmentation-reverse process. Lastly, it ensures the preservation of identical feature spaces for both original and augmented tokens, allowing for the seamless utilization of existing model architectures or pre-trained weights.
We assume token inputs instead of pixels for storage efficiency [36]. Thus, the input is a sequence of tokens that represents a pixel image . Since the tokens are stored in quantized form, each token means an embedding vector in a codebook. i.e., where for a codebook with size . The goal of TokenAdapt is to design an augmentation function that converts the input tokens to augmented tokens . Note that the augmented tokens should also be quantized .
Figure 4 depicts the overall pipeline of our TokenAdapt. When a token sequence and a pre-trained codebook is given, the token embedding corresponds to the codebook is obtained through lookup process, i.e., . The initial step of TokenAdapt is transforming into a representation in augmentation-compatible space . This transformation is achieved using a conversion function , i.e., . We apply augmentation at augmentation-compatible space and build an augmented representation . Then, the augmented representation is reverted to token embedding space by a reverse transformation , . At last, the augmented tokens in embedding space are quantized to codebook index as , where indicates the vector quantization process. In summary, the augmentation process of TokenAdapt is:
The conversion function and the reverse transformation are learned from paired data of tokens , where and denote tokens from an image and its augmented image , respectively. We use the cross-entropy loss between and as our objective function for learning our TokenAdapt module:
We particularly focus on geometric data augmentations (e.g., flip, resize, affine) and Mixup [58], which result in a performance drop for the baseline method, as shown in Table 1. Figure 3 shows the difference between direct augmentation to embedding space and augmentation with TokenAdapt . We decode the augmented tokens as images to help to understand differences in embedding space. The resulting images show that TokenAdapt effectively leverages existing pixel-based augmentations on tokens, thereby mitigating the image degradation caused by their direct application.
Once trained, TokenAdapt module is utilized to augment a sequence of tokens to an augmented sequence across various datasets and downstream tasks. Importantly, the augmentation-compatible feature space, not reliant on specific semantic knowledge or alignment to downstream tasks, renders it dataset-agnostic as well as task-agnostic. This inherent flexibility of our TokenAdapt module makes it highly generalizable, enabling direct application to other datasets without the need for additional training procedures.
3.3 ColorAdapt
As witnessed in Cubuk et al. [11], color-based augmentations are often more effective than geometric ones in certain datasets and play an important role in data augmentation strategies. For example, color-based augmentations are particularly useful in handling diverse environmental changes caused by varying lighting (e.g., day and night) and weather conditions (e.g., rain, fog, or snow). However, integrating them with TokenAdapt poses a challenge since most color-based augmentations are originally crafted to operate exclusively in the pixel domain.
To address this, we propose ColorAdapt, a new color-based token augmentation inspired by the method [24]. Formally, our ColorAdapt function is defined as:
where and represent the mean and standard deviation of each token embedding across its spatial dimension, computed for each channel.
Through transforming the global statistics within each channel of token embedding, ColorAdapt mimics color-changing augmentations in a pixel domain while preserving the encoded object structure. In Figure 5, we display the images decoded from various color-augmented tokens. This illustrates how ColorAdapt influences the visual characteristics of the augmented tokens. By thoughtfully incorporating ColorAdapt into our token augmentation strategy, we improve the data diversity, exploring new opportunities to train more robust and adaptable vision models.
| Method | Input | # of images |
Top 1 Acc. |
| Full pixels | Image | 1,281 k | 109 kB | 81.8 | 140.0 GB | ||||
| Uniform random sampling | 512 k | 107 kB | 74.0 | 54.6 GB (39%) | |||||
| C-score [26] based sampling | Image | 512 k | 104 kB | 73.3 | 53.3 GB (38%) | ||||
| Adjusting JPEG quality factor to 5 | 1,281 k | 9 kB | 74.6 | 11.0 GB (8%) | |||||
| Adjusting image resolution ( 0.2) | 1,281 k | 8 kB | 75.2 | 9.6 GB (7%) | |||||
| Ours | Token | 1,281 k | 1 kB | 75.5 | 1.4 GB (1%) |
3.4 Training procedure with TokenAdapt
We incorporate the trained TokenAdapt module to enable pixel-based data augmentation to tokens. Our evaluation focuses on two crucial computer vision tasks: (1) image classification and (2) semantic segmentation. For image classification, our goal is to enhance both model performance and robustness through the incorporation of our token augmentation strategies. Specifically, we stochastically apply pixel-based data augmentation to tokens using our TokenAdapt module. For color-based augmentations, we randomly choose between Emb-Noise [36] and our ColorAdapt, with a predetermined probability. By leveraging the diverse noise from Emb-Noise and retaining reliable object structure from our ColorAdapt, we observe an overall improvement in model performance.
In semantic segmentation, we leverage our TokenAdapt for pixel-based data augmentation as its application in image classification. However, a notable distinction exists: we do not employ mixup-based augmentations in semantic segmentation due to their limited performance benefits. This behavior aligns with recent studies [55, 46, 6] that exclude mixup-based augmentation in their training recipes. Additionally, we observe that token-level Random Resized Crop (RRC) does not enhance performance in segmentation due to the importance of structural integrity in pixel-level classification tasks. Thus, we apply geometric pixel-data augmentations using only our TokenAdapt.
4 Experiment
In this section, we validate the effectiveness of our proposed augmentation strategies across various scenarios where data augmentation is crucial. Initially, we evaluate our method’s performance in the context of storage-efficient ImageNet-1k classification. Furthermore, we demonstrate the adaptability of token-based vision model training to dense prediction tasks, showcasing additional improvements facilitated by our augmentation strategies. Additionally, we present experimental results illustrating how our strategies contribute to enhanced model generalizability in fine-grained classification and robustness evaluation. Lastly, we explore the applicability of our method to alternative input formats, demonstrating its extensibility. Unless otherwise stated, all the evaluations are performed on SeiT [36] framework.
4.1 Storage-efficient ImageNet
Table 2 summarizes the performance and compression ratio of storage-efficient ImageNet-1k classification. Our method is compared with image-based approaches, including uniform random sampling, C-score [26]-based sampling, adjusting image resolution, and adjusting JPEG quality factor. Despite a 1% compression ratio, our method achieves a 75.5% top-1 accuracy, outperforming various image-based approaches. In Table 3, our method consistently outperforms the token-based baseline by 1.5–4.1%p in top-1 accuracies. Notably, the effectiveness of our method becomes more pronounced as the available storage size decreases, emphasizing its efficiency in low-storage scenarios.
| Input |
Baseline | Ours | ||
| Token | 1.4 GB | 74.0 | 75.5 (+1.5) | ||
| 1.1 GB | 70.2 | 73.1 (+2.9) | |||
| 0.8 GB | 66.3 | 69.1 (+2.8) | |||
| 0.5 GB | 59.7 | 63.8 (+4.1) | |||
| 0.3 GB | 47.2 | 51.2 (+4.0) |
| Method | Flowers | Cars | iNat18 | iNat19 |
| Baseline | 93.5 | 79.7 | 43.1 | 50.1 |
| Ours | 93.6 | 83.9 | 47.5 | 55.5 |
| + | +0.1 | +4.2 | +4.4 | +5.4 |
| Method | Clean | Gauss. Noise | Gauss. Blur | ImageNet-R | Sketch | ObjectNet |
| Baseline | 74.0 | 50.7 | 62.6 | 25.5 | 22.6 | 15.7 |
| Ours | 75.5 (+1.5) | 58.6 (+7.9) | 66.8 (+4.2) | 30.2 (+4.7) | 27.7 (+5.1) | 18.3 (+2.6) |
4.2 Fine-grained classification
To further investigate the generalizability of our method, we conduct experiments on four different fine-grained datasets: Flowers [35], StandfordCars [27], iNat-18 [50], and iNat-19 [51]. Specifically, we tokenize each dataset and then fine-tune the token-trained model (ViT-B) on ImageNet-1k with the tokenized dataset. Since fine-grained datasets require classifying objects with intricate characteristics, they are particularly sensitive to undesirable noise (i.e., RRC in SeiT) or excessive perturbation (i.e., Emb-Noise) resulting from data augmentation in token embedding space. Table 4 shows the quantitative performance (top-1 accuracies) on fine-grained datasets. Our approach consistently outperforms the performance of the baseline across various datasets. These results confirm the effectiveness of our proposed method in fine-grained classification.
4.3 Evaluation on robustness datasets
We further evaluate the robustness of our method on a set of benchmarks, such as adding Gaussian noise [22] or Gaussian blur [22], ImageNet-R [23], Sketch [52], and ObjectNet [5]. The results in Table 5 demonstrate that our method consistently outperforms the baseline in terms of robustness accuracies. Previous studies [7, 47] observe that the key to the input pixel robustness depends on the pixel-level augmentations. Since our method enables pixel-based data augmentation to tokens, it achieves higher robustness accuracies than the baseline, which does not effectively exploit the pixel-based data augmentation.
| # of images | 127k | 76k | 25k | 13k |
| Storage | 138 MB | 83 MB | 28 MB | 14 MB |
| Baseline | 77.3 | 70.3 | 53.3 | 43.2 |
| Ours | 81.4 | 76.5 | 61.6 | 52.3 |
| + | +4.1 | +6.2 | +8.3 | +9.1 |
4.4 Classification with limited data
Data augmentation becomes more critical in limited data scenarios. To assess the impact of our approach in such situations, we conduct experiments on the ImageNet-100 benchmark by randomly sampling data based on varying data ratios. The results (top-1 accuracies), as depicted in Table 6, consistently demonstrate our method’s superiority over our baseline. Notably, the performance gap increases as the amount of training data decreases. With only 13k images, our method exhibits an improvement of approximately 21% over the baseline, indicating the effectiveness of our approach, especially in data-scarce scenarios.
4.5 ADE-20k semantic segmentation
A limitation of token-based learning is often substantiated by learning dense prediction tasks. We conjecture that our method could facilitate training for those tasks, such as segmentation tasks. We extract tokens from ADE-20k [62] and utilize the pre-trained ViT-B/16 model on tokenized ImageNet-1k to evaluate the effectiveness of our method in semantic segmentation. Subsequently, we fine-tune the pre-trained model using the tokenized ADE-20k dataset. Table 7 reports the mean intersection-over-union (mIoU) as the evaluation metric. The results demonstrate a notable performance increase (+2.5%p mIoU) with our data augmentation strategies compared to the baseline. These findings highlight the significance of addressing artifacts introduced by directly applying pixel-based data augmentation to token embeddings in semantic segmentation, a challenge effectively mitigated by our augmentation approach.
4.6 Analysis of TokenAdapt
We analyze the influence of our approach on training ViT-B for token-based ImageNet-1k classification, as depicted by the loss curves in Figure 6. Observing the training loss, our method consistently exhibits higher values than the baseline. Meanwhile, the test loss aligns with the trends of the baseline. This implies that the integration of token augmentation with our method effectively enhances data diversity, mitigating overfitting concerns. Notably, while the baseline’s test loss increases in later training epochs, our approach consistently decreases, highlighting its effectiveness in preventing the overfitting problem.
Furthermore, to demonstrate the efficacy of our strategies, we visualize the embedded feature representation in Figure 7. This illustration offers insight into the linear separability of augmented data points, particularly when pixel-based data augmentation is applied with our method. When pixel-based data augmentation is directly applied to token embedding without our method, there is significant overlap among augmented data points from different classes, indicating low linear separability. This overlap suggests that augmenting tokens without our method leads to substantial shifts in the training data distribution. Maintaining distributional similarity between clean and augmented data is crucial for model performance, as witnessed in Cubuk et al. [13], and significant distribution shifts can result in performance degradation. In contrast, our method shows high linear separability, demonstrating its effectiveness in diversifying the data while minimizing the distributional gap between clean and augmented data. Moreover, this behavior aligns with the observed trend in the case of applying data augmentation to images (DeiT-B [48] in Figure 7).
4.7 Ablation study
| Framework | ColorAdapt | TokenAdapt | Top 1 Acc. |
| SeiT [36] | ✘ | ✘ | 77.3 |
| ✔ | ✘ | 78.3 | |
| ✘ | ✔ | 80.4 | |
| ✔ | ✔ | 81.4 |
Impact of each augmentation strategy. In Table 8, we report the impact of proposed token augmentations. Notably, integrating ColorAdapt results in a 1.0%p performance boost, emphasizing the positive effect of preserving object structure during model training. Additionally, incorporating TokenAdapt yields an extra 3.1%p improvement. These results confirm that exploring token augmentations is a key factor in fully exploiting pre-extracted tokens.
| Input | Baseline | Ours |
| Token (ViT-VQGAN [56]) | 77.3 | 81.4 (+4.1) |
| Token (VQGAN [17]) | 81.8 | 83.9 (+2.1) |
| DCT coefficients | 69.5 | 71.0 (+1.5) |
Extensibility of our approach. In this section, we conjecture the broad applicability of our method across various scenarios. To explore its extensibility, we delve into two distinct situations. Firstly, we examine the integration of our method with an alternative tokenizer—the OpenImages-trained VQGAN tokenizer [17], alongside the ImageNet-1k-trained ViT-VQGAN tokenizer [56]. The results in Table 9 consistently demonstrate improved performance on ImageNet-100 across different tokenizers. Additionally, we investigate the adaptability of our approach to a different input format by focusing on Discrete Cosine Transform (DCT) coefficients. Inspired by JPEG compression, we apply DCT to each image patch, yielding quantized DCT coefficients utilzed as input data during model training. As shown in Table 9, our approach enhances top-1 accuracy, similar to the tokenizer scenario. This consistent improvement across diverse input formats underscores the universal applicability of our method, highlighting its potential as a widely applicable solution for diverse input representations.
4.8 Implementation details
We use the pre-trained ViT-VQGAN tokenizer [56] for token extraction. For our TokenAdapt module, we employ a single transformer block for conversion and reverse function. The TokenAdapt module is trained for one epoch on ImageNet-1k with a batch size of 128, a learning rate of 0.001 with cosine scheduling, and a weight decay of 0.05. We follow the training recipes from SeiT [36] and mmsegmentation [9] for token-based classification and semantic segmentation, respectively. Detailed hyperparameters and settings are available in the Appendix.
5 Conclusion
In this paper, we address a significant challenge in token-based vision model training: the limitations of integrating existing data augmentations designed for pixel images into token embedding due to input domain shift. To address this challenge, we propose simple yet effective token-based augmentation strategies, namely TokenAdapt and ColorAdapt. TokenAdapt realigns token embedding space to another embedding space compatible with pixel-based data augmentations. Once trained, the TokenAdapt module can be applied to different datasets and scenarios without the need for additional training. Additionally, we introduce ColorAdapt, a color-changing augmentation for tokens, to further diversify the training data. Our experimental results across diverse scenarios, including storage-efficient ImageNet-1k classification, fine-grained classification, robustness benchmarks, and ADE-20k semantic segmentation, consistently validate the effectiveness of our proposed method. Additionally, the consistent performance improvement across diverse input formats highlights the potential of our approach as a widely applicable solution for various input representations.
Appendix
In this supplementary material, we describe more details of our method. In Section A, we include more implementation details for our experiments (i.e., token-based image classification and semantic segmentation). Section B explores the data efficiency (Section B.1) and computational cost (Section B.2) of our TokenAdapt module. Additionally, Section C presents more qualitative results showcasing the effectiveness of our proposed TokenAdapt and ColorAdapt for augmenting tokens.
Appendix A Implementation details
Token-based image classification. For token ImageNet-1k training, we follow the training recipe from SeiT [36], adjusting only the warm-up epochs for stable convergence. We conducted token-based image classification on the ViT-B/16 model [16, 48] and used a learning rate of 0.0015 with cosine scheduling and a weight decay of 0.1. The model was trained for 300 epochs with a batch size of 1024. Regarding data augmentations, SeiT [36] incorporates Token-RRC, Token-EDA, Token-CutMix, and Emb-Noise. As a default, we integrate these token-based data augmentation methods and further enhance token diversity by applying additional augmentations using our TokenAdapt and ColorAdapt. Specifically, we apply pixel-based data augmentations (e.g., RRC, hFlip, affine transformations, Mixup, and CutMix) to tokens using our TokenAdapt module with a probability of 0.5. We adopt the hyperparameters for data augmentation proposed in DeiT [48]. For token-based fine-grained classification, following the SeiT, we used the training recipe of DeiT. In all experiments, when the number of data points decreased, we modified the number of total training iterations to ensure a fair comparison.
Token-based semantic segmentation. For the token ADE-20k dataset preparation, we initially resize the entire ADE-20k dataset to . Following the procedure described in SeiT, we utilize the ViT-VQGAN tokenizer [56] to extract tokens from the resized images. Token-based semantic segmentation on ADE-20k follows the training recipe from mmsegmentation [9], using the DeiT-B/16 model with UperNet [54]. The training involves a learning rate of 6e-5 with polynomial scheduling and a weight decay of 0.01. The model was trained for 80k iterations with a batch size of 16. Regarding data augmentations, as mentioned in Section 3.2, we do not employ mixup-based augmentations in token-based semantic segmentation. Furthermore, we observe that token-level Random Resized Crop (RRC) does not enhance performance due to the importance of maintaining structural integrity in pixel-level classification tasks. Thus, we apply geometric pixel-based data augmentations only using our TokenAdapt module.
| # of training data | 100 | 1k | 5k | 256k | 1,281k |
| Top 1 Acc. | 75.1 | 75.4 | 75.4 | 75.4 | 75.5 |
| (IN-1k CLS.) |
| Method | #params | GPU hours | Top 1 Acc. |
| Baseline | 22M | 5.9 | 77.3 |
| Baseline | 22M+172M | 49.2 (+734%) | 80.1 (+2.8) |
| Ours | 22M+3.7M | 6.4 (+8%) | 81.4 (+4.1) |
Appendix B Efficiency analysis of TokenAdapt
B.1 Data efficiency of TokenAdapt
To investigate the data efficiency of the TokenAdapt module, we explore the impact of reduced training data for the TokenAdapt module training. Specifically, We vary the number of training data for the TokenAdapt module training and subsequently train a storage-efficient ImageNet-1k classification model using the trained TokenAdapt module. Table A shows that even with a reduced number of training data for our TokenAdapt module, the performance degradation is minimal. Notably, the top-1 accuracy remains more than 1%p higher than the baseline (74.0% top-1 accuracy) even with only 100 training data. This robust performance suggests that our TokenAdapt module is not significantly affected by variations in the training data size.
B.2 Computational costs
To validate the computational efficiency of TokenAdapt, we compare ViT-S models trained on ImageNet-100 with different training strategies (Table B.2). Three scenarios are considered: (1) only token-based augmentations (baseline), (2) pixel-based augmentations in decoded images (baseline), and (3) pixel-based augmentations using our TokenAdapt module. For baseline, utilizing the ViT-VQGAN decoder (86M) and encoder (86M) during each forward computation significantly increases training time by over 8 times compared to the baseline, making it impractical for addressing the data augmentation challenge. In contrast, our TokenAdapt module, requiring only 3.7M parameters, increases training time by only 8% compared to the baseline while achieving a remarkable performance improvement over 3%p in top-1 accuracy. Notably, TokenAdapt even outperforms baseline, indicating that the full decoding-augmentation-encoding process may introduce undesirable noise during tokenization. These results demonstrate the efficient handling of the data augmentation challenge by our TokenAdapt module. We note that the TokenAdapt module can be removed during the inference phase.
Appendix C More qualitative examples
We demonstrate the effectiveness of the proposed TokenAdapt and ColorAdapt by decoding the augmented tokens into images using the ViT-VQGAN decoder [56]. In Figure A, we compare the direct application of pixel-based data augmentations to token embedding with augmentations using our TokenAdapt module. As mentioned in Section 3.1, direct application of hFlip or affine transformations causes disruption in both the diagonal line and the object’s silhouette due to spatial information collapse. Additionally, data augmentations related to interpolation (e.g., Resize or Mixup) can result in undesired artifacts. Mixup, in particular, occasionally causes significant disruption of objects, leading to substantial performance degradation as shown in Table 1. The occurrence of these unexpected artifacts makes it difficult to use the data augmentations, which are widely used in existing pixel image domains. In contrast, The tokens that are augmented using our TokenAdapt module exhibit more reasonable results, mitigating the image degradation caused by the direct application of pixel-based augmentations to tokens. Consequently, TokenAdapt consistently improves model performance across various scenarios by effectively leveraging pixel-based data augmentation in a token domain.
In Figure B, we present the images decoded from color-augmented tokens by various existing color-based data augmentations. For pixel-based color augmentation, we employ brightness and contrast, which are widely adopted for pixel-based vision model training. Specifically, we used brightness and contrast functions following the implementation [38]. For token-based color augmentation, we employ Emb-Noise [36] with the same optimized hyperparameters used in token-based vision model training [36]. Figure B illustrates the impact of the proposed ColorAdapt on the visual characteristics of augmented tokens. Unlike the existing color-based augmentations, our ColorAdapt effectively preserves object structure while introducing significant color variations. Recognizing the importance of maintaining object structure in diverse vision tasks (e.g., fine-grained classification, semantic segmentation), our ColorAdapt opens up new possibilities for training more robust and adaptable vision models.